# Occupied Processes: Going with the Flow

### v4

@@ 3.3 Feynman-Kac’s Formula for the Backward Heat Equation
$v(\mathfrak{o},x)=\operatorname{\mathbb{E}}_{\mathfrak{o},x}^{\operatorname{\mathbb{Q}}}[\varphi(\operatorname{\mathcal{O}}_{T},X_{T})]\mathrel{\mathop{\ordinarycolon}}=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\varphi(\mathfrak{o}+\operatorname{\mathcal{O}}_{T-t}^{x},x+X_{T-t})],\quad\mathfrak{o}(\operatorname{\mathbb{R}})=t,$ (36)
@@
$\textnormal{v}(t,\omega)=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[L_{T}^{0}|\operatorname{\mathcal{F}}_{t}](\omega)=L^{0}_{t}(\omega)+\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}\big{[}L_{T-t}^{|x|}\big{]}\big{|}_{x=\omega_{t}},\quad\omega\in\Omega_{T},$ (37)
is not continuous in $\omega$ (with respect to the supremum norm), regardless of the definition of $L^{0}_{t}\mathrel{\mathop{\ordinarycolon}}\Omega_{T}\to\operatorname{\mathbb{R}}$. First, one may set $L^{0}_{t}(\omega)=\lim_{\varepsilon\to 0}\frac{1}{2\varepsilon}\int_{0}^{t}\mathds{1}_{B_{\varepsilon}}(\omega_{s})ds$ as the occupation measures $\operatorname{\mathcal{O}},\tilde{\operatorname{\mathcal{O}}}$ almost surely coincide under the Wiener measure. However, if $(\omega^{N})$ is a sequence of piecewise linear paths that converges to $\omega$ (e.g. using Schauder’s theorem) it may happen that $L^{0}_{t}(\omega^{N})=\infty$ as the occupation measure of $\omega^{N}$ is not absolutely continuous with respect to $\lambda$. Another possibility is $L^{0}_{t}(\omega)\mathrel{\mathop{\ordinarycolon}}=\lim_{\varepsilon\to 0}\frac{1}{2\varepsilon}\int_{0}^{t}\mathds{1}_{B_{\varepsilon}}(X_{s})d\langle X\rangle_{s}(\omega)$ when $\langle X\rangle$ is well-defined. But $L^{0}_{t}(\omega^{N})=0$ for the above sequence $(\omega^{N})$, leading again to the discontinuity of $L^{0}_{t}$ with respect to $\omega$. Finally, we could derive the local time at zero (or any other level) from Tanaka’s formula [74, Chapter VI]. However, the expression would involve stochastic integrals which are infamously discontinuous with respect to the input path. This representation is thus subject to the same fate as the previous ones.
@@ 3.4 Occupied SDEs and Dirichlet Problems
$|b(\mathfrak{o},x)|^{2}$ $\leq K^{2}[1+\varrho(\mathfrak{o},x)^{2}],\quad|\sigma(\mathfrak{o},x)|^{2}\leq K^{2}[1+\varrho(\mathfrak{o},x)],$ (41)
$|f(\mathfrak{o}^{\prime},x^{\prime})-f(\mathfrak{o},x)|$ $\leq K\varrho(\mathfrak{o}^{\prime}-\mathfrak{o},x^{\prime}-x),\quad f\in\{b,\sigma\}.$ (42)
@@
Notice the unusual linear growth condition for the squared diffusion coefficient $\sigma^{2}$ in (41) instead of $\sigma$ as in the classical case [74, Chapter IX]. This is because $\sigma^{2}(\operatorname{\mathcal{O}},X)$ appears in the dynamics of the occupation flow.
@@
$v(\mathfrak{o},x)=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\varphi(\operatorname{\mathcal{O}}_{\tau_{\Theta}}^{\mathfrak{o},x},X_{\tau_{\Theta}}^{\mathfrak{o},x})],\quad(\mathfrak{o},x)\in\overline{\operatorname{\mathcal{D}}_{\Theta}}.$ (46)
@@
$v(\mathfrak{o},x)=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\varphi(\tilde{\operatorname{\mathcal{O}}}_{T}^{\mathfrak{o},x},X_{T}^{\mathfrak{o},x})],\quad(\mathfrak{o},x)\in\overline{\operatorname{\mathcal{D}}_{T}}.$ (49)
@@ 4 Unified Markovian Lift of Financial Instruments
$\xi(\omega)=\varphi(F_{T},X_{T})(\omega),\quad\varphi\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathfrak{F}}\times\operatorname{\mathbb{R}}\to\operatorname{\mathbb{R}},\quad\ \omega\in\Omega,$ (50)
while $(F,X)$ is a Markov process. The feature space $\operatorname{\mathfrak{F}}$ is chosen to be ”nicer” than $\Omega$, e.g., a finite-dimensional Euclidean space $\operatorname{\mathfrak{F}}=\operatorname{\mathbb{R}}^{d}$ or $\operatorname{\mathfrak{F}}=\operatorname{\mathcal{M}}$ in the case of occupation flows. Due to the Markov procperty, the derivatives price will be function of these variables as well, leading to tractable numerical methods to solve the corresponding pricing PDE in the projected space. Setting $F=\tilde{\operatorname{\mathcal{O}}}$ or $F=\operatorname{\mathcal{O}}$ then leads to occupied pricing PDEs, allowing the underlying to include the occupation flow in its dynamics as well. The use of occupied SDEs in financial modeling will be discussed in Sections 5 and 6. It turns out that the relevant features in finance are very often expressible in terms of the occupation flow, either with respect to calendar time and variance time as shown next.
@@ 4.1 Exotic Options and Structured Products
$\varphi(\tilde{\operatorname{\mathcal{O}}}_{T},X_{T})=\phi(\tilde{\operatorname{\mathcal{O}}}_{T}(A),X_{T}),\quad\phi\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}^{2}\to\operatorname{\mathbb{R}},\quad A\in\operatorname{\mathcal{B}}(\operatorname{\mathbb{R}}).$ (51)
@@ 4.1.1 Deep BSDE Solver
$\inf_{\pi,\theta}\ \operatorname{\mathcal{L}}(\pi,\theta),\quad\operatorname{\mathcal{L}}(\pi,\theta)\mathrel{\mathop{\ordinarycolon}}=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[|\xi-Y_{T}^{\pi,\theta}|^{2}].$ (55)
@@ 4.2 Variance Derivatives
$\text{P\&L}=\frac{1}{T}\int_{0}^{T}\mathds{1}_{[x_{1},x_{2}]}(X_{t})V_{t}dt\;\ -\;\ K^{2}_{x_{1},x_{2}}\ ,\qquad x_{1}<x_{2}.$ (56)
@@ 5 Occupied Volatility
ensuring that vanilla call and put options are correctly priced; see, e.g., [20]. On the other hand, occupied volatility is a particular example of path-dependent volatility [20, 52, 53, 56, 59] and thus enjoys similar benefits. Notably, as the occupation flow is endogenous, the model remains complete, i.e., market makers can replicate any contingent claim $\xi$ by dynamically trading the underlying $X$.
@@
$d\mathcal{O}^{\kappa}_{t}=\delta_{X_{t}}e^{\kappa t}dt,\quad\kappa\geq 0.$ (63)
@@
Importantly, $(\operatorname{\mathcal{O}}^{\kappa},X)$ is still a Markov process. The exponential weight in $\operatorname{\mathcal{O}}^{\kappa}$ puts more emphasis on recent events. In particular, the barycenter of $\mathcal{O}^{\kappa}$ coincide with the exponentially weighted moving average (EWMA) of the asset, namely
$\frac{1}{\mathcal{O}^{\kappa}_{t}(\mathbb{R})}\int x\mathcal{O}^{\kappa}_{t}(dx)=\frac{1}{1-e^{-\kappa t}}\int_{0}^{t}X_{u}\kappa e^{-\kappa(t-u)}du.$ (65)
The calendar time occupation flow is recovered by setting $\kappa=0$, so all the subsequent results will presented in terms of the more general object $\operatorname{\mathcal{O}}^{\kappa}$. Note that calendar time may still preferred in the applications due to the unified framework it provides for the pricing of exotic options as in Fig. 3.
@@
$\sigma(\operatorname{\mathcal{O}}_{t}^{\kappa},X_{t})=\Sigma(\Upsilon(\operatorname{\mathcal{O}}_{t}^{\kappa},X_{t})),\quad\Sigma(y)=-\frac{\alpha}{\beta}+\gamma y^{-\beta},$ (66)
with parameters $\alpha,\beta,\gamma>0$ and trend functional,
$\Upsilon(\operatorname{\mathcal{O}}_{t}^{\kappa},X_{t})=\frac{X_{t}}{\frac{1}{\operatorname{\mathcal{O}}_{t}^{\kappa}(\operatorname{\mathbb{R}})}\int_{\operatorname{\mathbb{R}}}x\operatorname{\mathcal{O}}_{t}^{\kappa}(dx)}=\frac{X_{t}}{\frac{1}{1-e^{-\kappa t}}\int_{0}^{t}X_{u}\kappa e^{-\kappa(t-u)}du}.$ (67)
@@
Figure 6: Guyon’s toy volatility model with $\sigma$ given in (66). One simulations using the Euler-Maruyama scheme in 5.1. Parameters: $x_{0}=100,T=1,\kappa=12$, and $(\alpha,\beta,\gamma)=(2.1,1.2,1.9)$.
@@ 5.1 Simulation of Occupied SDEs
$C_{m}=[x_{m}-\varepsilon_{m-1},x_{m}+\varepsilon_{m}),\quad\varepsilon_{m}=\frac{x_{m+1}-x_{m}}{2},\quad(\varepsilon_{0}=x_{1},\varepsilon_{M}=\infty).$
@@
$\frac{dX_{t}}{X_{t}}=(r-q)dt+\sigma^{M}(O_{t}^{\kappa},X_{t})dW_{t},\qquad dO_{t,m}^{\kappa}=\mathds{1}_{C_{m}}(X_{t})e^{\kappa t}dt,$ (68)
with $O_{t}^{\kappa}=(\operatorname{\mathcal{O}}_{t}^{\kappa}(C_{1}),\ldots,\operatorname{\mathcal{O}}_{t}^{\kappa}(C_{M}))\in\operatorname{\mathbb{R}}^{M}$. The projected volatility function $\sigma^{M}\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}^{M+1}\to\operatorname{\mathbb{R}}$ is given by $\sigma^{M}(O_{t}^{\kappa},X_{t})=\sigma(\sum_{m=1}^{M}O_{t,m}^{\kappa}\delta_{x_{m}},X_{t})$, which converges to $\sigma(\operatorname{\mathcal{O}}_{t}^{\kappa},X_{t})$ provided that $\sigma(\mathfrak{o},x)$ is weakly continuous in $\mathfrak{o}$. We can then simulate $X$ using standard numerical schemes for SDEs [65, 51]. To wit, consider a regular time grid $t_{n}=n\delta t$, $\delta t=\frac{T}{N}$, $N\in\operatorname{\mathbb{N}}$. Given $X_{0}\in\operatorname{\mathbb{R}}_{+}$ and $O^{\text{\tiny tmp}}=0\in\operatorname{\mathbb{R}}^{M}$, we then generate sample paths from (68) as follows,
$\begin{cases}O^{\text{\tiny tmp}}_{m}&\leftarrow\ O^{\text{\tiny tmp}}_{m}+e^{\kappa t_{n}}\delta t,\quad X_{t_{n}}\ \in\ C_{m},\\[1.7pt] \sigma_{t_{n}}&=\ \sigma^{M}(O^{\text{\tiny tmp}},X_{t_{n}}),\\[1.7pt] X_{t_{n+1}}&=\ X_{t_{n}}\exp\{\sigma_{t_{n}}\sqrt{\delta t}Z_{n}+(r-q-\frac{1}{2}\sigma_{t_{n}}^{2})\delta t\},\end{cases}$
@@
$\frac{dX_{t}}{X_{t}}=(r-q)dt+\sigma((X_{s})_{s\leq t})dW_{t},$
typically involves discretization of the form $\sigma((X_{s})_{s\leq t_{n}})\approx\sigma^{N}_{n}(X_{t_{1}},\ldots,X_{t_{n}})$ for some function $\sigma^{N}_{n}\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}^{n}\to\operatorname{\mathbb{R}}$. Note the input dimension of $\sigma^{N}_{n}$ grows linearly with $n$, which becomes highly problematic when the total number of time steps $N$ is large. Using OSDEs, however, we can decouple the dimensionality of the space and time discretization, which allows for efficient simulations as shown in Example 13 and Section 5.3.

### v5

@@ 3.3 Feynman-Kac’s Formula for the Backward Heat Equation
$v(\mathfrak{o},x)=\operatorname{\mathbb{E}}_{\mathfrak{o},x}^{\operatorname{\mathbb{Q}}}[\varphi(\operatorname{\mathcal{O}}_{T},X_{T})]\mathrel{\mathop{\ordinarycolon}}=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\varphi(\mathfrak{o}+\operatorname{\mathcal{O}}_{T-t}^{x},x+X_{T-t})],\qquad\mathfrak{o}(\operatorname{\mathbb{R}})=t,$ (36)
@@
$\textnormal{v}(t,\omega)=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[L_{T}^{0}|\operatorname{\mathcal{F}}_{t}](\omega)=L^{0}_{t}(\omega)+\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}\big[L_{T-t}^{|x|}\big]\big|_{x=\omega_{t}},\qquad\omega\in\Omega_{T},$ (37)
is not continuous in $\omega$ (with respect to the supremum norm), regardless of the definition of $L^{0}_{t}\mathrel{\mathop{\ordinarycolon}}\Omega_{T}\to\operatorname{\mathbb{R}}$. First, one may set $L^{0}_{t}(\omega)=\lim_{\varepsilon\to 0}\frac{1}{2\varepsilon}\int_{0}^{t}\mathds{1}_{B_{\varepsilon}}(\omega_{s})ds$ as the occupation measures $\operatorname{\mathcal{O}},\tilde{\operatorname{\mathcal{O}}}$ almost surely coincide under the Wiener measure. However, if $(\omega^{N})$ is a sequence of piecewise linear paths that converges to $\omega$ (e.g. using Schauder’s theorem) it may happen that $L^{0}_{t}(\omega^{N})=\infty$ as the occupation measure of $\omega^{N}$ is not absolutely continuous with respect to $\lambda$. Another possibility is $L^{0}_{t}(\omega)\mathrel{\mathop{\ordinarycolon}}=\lim_{\varepsilon\to 0}\frac{1}{2\varepsilon}\int_{0}^{t}\mathds{1}_{B_{\varepsilon}}(X_{s})d\langle X\rangle_{s}(\omega)$ when $\langle X\rangle$ is well-defined. But $L^{0}_{t}(\omega^{N})=0$ for the above sequence $(\omega^{N})$, leading again to the discontinuity of $L^{0}_{t}$ with respect to $\omega$. Finally, we could derive the local time at zero (or any other level) from Tanaka’s formula [75, Chapter VI]. However, the expression would involve stochastic integrals which are infamously discontinuous with respect to the input path. This representation is thus subject to the same fate as the previous ones.
@@ 3.4 Occupied SDEs and Dirichlet Problems
$|b(\mathfrak{o},x)|^{2}$ $\leq K^{2}[1+\varrho(\mathfrak{o},x)^{2}],\qquad|\sigma(\mathfrak{o},x)|^{2}\leq K^{2}[1+\varrho(\mathfrak{o},x)],$ (41)
$|f(\mathfrak{o}^{\prime},x^{\prime})-f(\mathfrak{o},x)|$ $\leq K\varrho(\mathfrak{o}^{\prime}-\mathfrak{o},x^{\prime}-x),\qquad f\in\{b,\sigma\}.$ (42)
@@
Notice the unusual linear growth condition for the squared diffusion coefficient $\sigma^{2}$ in (41) instead of $\sigma$ as in the classical case [75, Chapter IX]. This is because $\sigma^{2}(\operatorname{\mathcal{O}},X)$ appears in the dynamics of the occupation flow.
@@
$v(\mathfrak{o},x)=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\varphi(\operatorname{\mathcal{O}}_{\tau_{\Theta}}^{\mathfrak{o},x},X_{\tau_{\Theta}}^{\mathfrak{o},x})],\qquad(\mathfrak{o},x)\in\overline{\operatorname{\mathcal{D}}_{\Theta}}.$ (46)
@@
$v(\mathfrak{o},x)=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\varphi(\tilde{\operatorname{\mathcal{O}}}_{T}^{\mathfrak{o},x},X_{T}^{\mathfrak{o},x})],\qquad(\mathfrak{o},x)\in\overline{\operatorname{\mathcal{D}}_{T}}.$ (49)
@@ 4 Unified Markovian Lift of Financial Instruments
$\xi(\omega)=\varphi(F_{T},X_{T})(\omega),\qquad\varphi\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathfrak{F}}\times\operatorname{\mathbb{R}}\to\operatorname{\mathbb{R}},\qquad\ \omega\in\Omega,$ (50)
while $(F,X)$ is a Markov process. The feature space $\operatorname{\mathfrak{F}}$ is chosen to be ”nicer” than $\Omega$, e.g., a finite-dimensional Euclidean space $\operatorname{\mathfrak{F}}=\operatorname{\mathbb{R}}^{d}$ or $\operatorname{\mathfrak{F}}=\operatorname{\mathcal{M}}$ in the case of occupation flows. Due to the Markov property, the derivatives price will be function of these variables as well, leading to tractable numerical methods to solve the corresponding pricing PDE in the projected space. When $\operatorname{\mathfrak{F}}=\operatorname{\mathcal{M}}$, we stress that the occupation flow needs to be discretized in the implementation, leading to a finite-dimensional Markovian lift given by finitely many occupation times. Setting $F=\tilde{\operatorname{\mathcal{O}}}$ or $F=\operatorname{\mathcal{O}}$ then leads to occupied pricing PDEs, allowing the underlying to include the occupation flow in its dynamics as well. The use of occupied SDEs in financial modeling will be discussed in Sections 5 and 6. It turns out that the relevant features in finance are very often expressible in terms of the occupation flow, either with respect to calendar time and variance time as shown next.
@@ 4.1 Exotic Options and Structured Products
$\varphi(\tilde{\operatorname{\mathcal{O}}}_{T},X_{T})=\phi(\tilde{\operatorname{\mathcal{O}}}_{T}(A),X_{T}),\qquad\phi\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}^{2}\to\operatorname{\mathbb{R}},\qquad A\in\operatorname{\mathcal{B}}(\operatorname{\mathbb{R}}).$ (51)
@@ 4.1.1 Deep BSDE Solver
$\inf_{\pi,\theta}\ \operatorname{\mathcal{L}}(\pi,\theta),\qquad\operatorname{\mathcal{L}}(\pi,\theta)\mathrel{\mathop{\ordinarycolon}}=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[|\xi-Y_{T}^{\pi,\theta}|^{2}].$ (55)
@@ 4.2 Variance Derivatives
$\text{P\&L}=\frac{1}{T}\int_{0}^{T}\mathds{1}_{[x_{1},x_{2}]}(X_{t})V_{t}dt\;\ -\;\ K^{2}_{x_{1},x_{2}}\ ,\hskip 17.00024ptx_{1}<x_{2}.$ (56)
@@ 5 Occupied Volatility
ensuring that vanilla call and put options are correctly priced; see, e.g., [20]. On the other hand, occupied volatility is a particular example of path-dependent volatility [20, 52, 53, 56, 59] and thus enjoys similar benefits. Notably, as the occupation flow is endogenous, the model remains complete, i.e., every contingent claim $\xi$ can be perfectly replicated by dynamically trading the underlying $X$.
@@
$d\mathcal{O}^{\kappa}_{t}=\delta_{X_{t}}e^{\kappa t}dt,\qquad\kappa\geq 0.$ (63)
[…]
@@
By increasing the exponential parameter $\kappa$, the exponential time occupation flow puts more emphasis on the present, thus making volatility more reactive to recent shocks. Crucially, $(\operatorname{\mathcal{O}}^{\kappa},X)$ remains a Markov process when $\kappa>0$, enabling efficient simulation of (64).
@@
$\sigma(\operatorname{\mathcal{O}}_{t}^{\kappa},X_{t})=\Sigma(\Upsilon(\operatorname{\mathcal{O}}_{t}^{\kappa},X_{t})),\qquad\Sigma(y)=-\frac{\alpha}{\beta}+\gamma y^{-\beta},$ (65)
with parameters $\alpha,\beta,\gamma>0$, and functional
$\Upsilon(\operatorname{\mathcal{O}}_{t}^{\kappa},X_{t})=\frac{X_{t}}{\frac{1}{\operatorname{\mathcal{O}}_{t}^{\kappa}(\operatorname{\mathbb{R}})}\int_{\operatorname{\mathbb{R}}}x\operatorname{\mathcal{O}}_{t}^{\kappa}(dx)}.$ (66)
The ratio (66) characterizes the trend of the asset by noting that the barycenter of $\mathcal{O}^{\kappa}$ coincide with the exponentially weighted moving average (EMA) of the asset, namely
$\text{EMA}_{t}=\frac{1}{1-e^{-\kappa t}}\int_{0}^{t}X_{u}\kappa e^{-\kappa(t-u)}du=\frac{1}{\mathcal{O}^{\kappa}_{t}(\mathbb{R})}\int x\mathcal{O}^{\kappa}_{t}(dx).$ (67)
@@
Figure 6: Guyon’s toy volatility model with $\sigma$ given in (65). One simulations using the Euler-Maruyama scheme in Section 5.1. Parameters: $x_{0}=100,T=1,\kappa=12$, and $(\alpha,\beta,\gamma)=(2.1,1.2,1.9)$.
@@ 5.1 Simulation of Occupied SDEs
$C_{m}=[x_{m}-\varepsilon_{m-1},x_{m}+\varepsilon_{m}),\qquad\varepsilon_{m}=\frac{x_{m+1}-x_{m}}{2},\qquad(\varepsilon_{0}=x_{1},\varepsilon_{M}=\infty).$
@@
$\frac{dX_{t}}{X_{t}}=(r-q)dt+\sigma^{M}(O_{t}^{\kappa},X_{t})dW_{t},\hskip 17.00024ptdO_{t,m}^{\kappa}=\mathds{1}_{C_{m}}(X_{t})e^{\kappa t}dt,$ (68)
with $O_{t}^{\kappa}=(\operatorname{\mathcal{O}}_{t}^{\kappa}(C_{1}),\ldots,\operatorname{\mathcal{O}}_{t}^{\kappa}(C_{M}))\in\operatorname{\mathbb{R}}^{M}$. In light of (68), $(O^{\kappa},X)$ is an $M+1$--dimensional Markov process. Moreover, the projected volatility
$\sigma^{M}\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}^{M+1}\to\operatorname{\mathbb{R}},\hskip 17.00024pt\sigma^{M}(O_{t}^{\kappa},X_{t})=\sigma\Big(\sum\limits_{m=1}^{M}O_{t,m}^{\kappa}\delta_{x_{m}},X_{t}\Big),$
converges to $\sigma(\operatorname{\mathcal{O}}_{t}^{\kappa},X_{t})$ when $\sigma(\mathfrak{o},x)$ is weakly continuous in $\mathfrak{o}$. We can then simulate $X$ using standard numerical schemes for SDEs [65, 51]. To wit, consider a regular time grid $t_{n}=n\delta t$, $\delta t=\frac{T}{N}$, $N\in\operatorname{\mathbb{N}}$. Given $X_{0}\in\operatorname{\mathbb{R}}_{+}$ and $O^{\text{\tiny tmp}}=0\in\operatorname{\mathbb{R}}^{M}$, we then generate sample paths from (68) as follows,
$\begin{cases}O^{\text{\tiny tmp}}_{m}&\leftarrow\ O^{\text{\tiny tmp}}_{m}+e^{\kappa t_{n}}\delta t,\qquad X_{t_{n}}\ \in\ C_{m},\\[1.7pt] \sigma_{t_{n}}&=\ \sigma^{M}(O^{\text{\tiny tmp}},X_{t_{n}}),\\[1.7pt] X_{t_{n+1}}&=\ X_{t_{n}}\exp\{\sigma_{t_{n}}\sqrt{\delta t}Z_{n}+(r-q-\frac{1}{2}\sigma_{t_{n}}^{2})\delta t\},\end{cases}$
@@
$\frac{dX_{t}}{X_{t}}=(r-q)dt+\sigma(t,\omega)dW_{t},$
typically involves discretization of the form $\sigma(t_{n},\omega)\approx\sigma^{N}_{n}(X_{t_{1}}(\omega),\ldots,X_{t_{n}}(\omega))$ for some function $\sigma^{N}_{n}\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}^{n}\to\operatorname{\mathbb{R}}$. Note the input dimension of $\sigma^{N}_{n}$ grows linearly with $n$, which becomes highly problematic when the total number of time steps $N$ is large. In contrast, occupied SDEs decouple the dimensionality of the space and time discretization, allowing for efficient simulations as shown in Example 13 and Section 5.3.